At 2:00 p.m. two cars start toward each other from towns 240 miles apart. If the rate of one car is 10 mph faster than the other, how fast does each car go if they meet at 5:00 p.m.?
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(9135)

(x+10)*3+3*x=240
x=35
x+10=45
check
3*35+3*45=240
105+135=240
ok

Answer by stanbon(57282)

At 2:00 p.m. two cars start toward each other from towns 240 miles apart. If the rate of one car is 10 mph faster than the other, how fast does each car go if they meet at 5:00 p.m.?
--------------------
Slower Car DATA:
distance = x ; time = 3 hrs ; rate = d/t = x/3 hrs
-------------------------
Faster Car DATA:
distance = (240-x) ; time = 3 hrs ; rate = d/t = (240-x)/3
-----------------------------
Equation:
faster rate - slower rate = 10 mph
(240-x)/3 - x/3 = 10 mph
240-x - x = 30
-2x = -210
x = 105 miles
----
slower car rate = 105/3 = 35 mph
faster car rate = 35+10 = 45 mph
===================================
Cheers,
Stan H.